In a particular city, pulse rates are normally distributed, with a mean of 70.4 beats per minute and a standard deviation of 9.9 beats per minute. A researcher is doing a study that requires volunteers whose pulse rates are in the highest 30%. What is the pulse rate that separates the highest 30% from the rest? Include the correct units!
Answer #1

Solution:-

Given that,

mean = \mu = 70.4

standard deviation = \sigma = 9.9

Using standard normal table,

P(Z > z) = 30%

= 1 - P(Z < z) = 0.30  

= P(Z < z) = 1 - 0.30

= P(Z < z ) = 0.70

= P(Z < 0.52) = 0.70  

z = 0.52

Using z-score formula,

x = z * \sigma + \mu

x = 0.52 * 9.9 + 70.4

x = 75.548